OpenStudy (anonymous):

If f '(x)> 0 for all x in (-∞,∞), then we know that: a. f(x) can have no relative extrema. b. f(x) can have no inflection points. c. For some x-value, f(x) > 0. d. None of the above are true.

OpenStudy (mathmagician):

b

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

Disagree with mathmagician. Here's my counterexample: f'' = 2x f' = x^2 +1 f = (1/3)x^3 + x f' Is always positive, but there's an inflection point at x=0.

OpenStudy (anonymous):

However, consider what happens to the derivative of a function at its relative extrema.

OpenStudy (anonymous):

a

OpenStudy (anonymous):

since we'll never have f'(x)=0, there can be no critical points such that f'(x)=0, so there will be no extremes.

OpenStudy (anonymous):

And as a counter example for c, consider: